# EchoVideo: Identity-Preserving Human Video Generation by Multimodal Feature Fusion

### v1

@@ Abstract
Recent advancements in video generation have significantly impacted various downstream applications, particularly in identity-preserving video generation (IPT2V). However, existing methods struggle with "copy-paste" artifacts and low similarity issues, primarily due to their reliance on low-level facial image information. This dependence can result in rigid facial appearances and artifacts reflecting irrelevant details. To address these challenges, we propose EchoVideo, which employs two key strategies: (1) an Identity Image-Text Fusion Module (IITF) that integrates high-level semantic features from text, capturing clean facial identity representations while discarding occlusions, poses, and lighting variations to avoid the introduction of artifacts; (2) a two-stage training strategy, incorporating a stochastic method in the second phase to randomly utilize shallow facial information. The objective is to balance the enhancements in fidelity provided by shallow features while mitigating excessive reliance on them. This strategy encourages the model to utilize high-level features during training, ultimately fostering a more robust representation of facial identities. EchoVideo effectively preserves facial identities and maintains full-body integrity. Extensive experiments demonstrate that it achieves excellent results in generating high-quality, controllability and fidelity videos.
@@ 2.2 Identity-preserving Video Generation
Identity-preserving generation aims to retain distinct identity attributes in generated images or videos. Building on the remarkable success of diffusion models in text-to-image synthesis, numerous methods [42, 43, 44, 45, 46] have emerged that leverage these models for identity-preserving image generation. However, achieving identity preservation at the video level introduces additional complexity, as it requires maintaining identity consistency across frames while ensuring the naturalness of facial movements. Among the U-Net-based approaches, MagicMe [7] relies on fine-tuning with reference images, while ID-Animator [8] introduces a face adapter to encode facial features, allowing for tune-free operation. However, both methods have limitations regarding facial similarity and overall video quality. In contrast, several open-source DiT-based methods, such as ConsisID [10] and Magic Mirror [9], leverage the robust CogVideoX [37] model as a foundation, integrating facial information through cross-attention mechanisms to achieve identity preservation. Validation of the ConsisID [10] model reveals that, despite generating high-quality videos and maintaining strong facial similarity, it suffers from significant "copy-paste" artifacts. This issuearises from the model’s structural design, which fails to adequately decouple facial features. To address these challenges, we propose EchoVideo, which incorporates a multimodal feature fusion module named IITF. This module captures only the high-level semantic and identity information of the face while disregarding irrelevant factors such as occlusions, poses, and lighting variations. Additionally, the fused semantic features help resolve semantic conflicts across different modalities, resulting in the generation of more stable and coherent character representations in videos.
@@ 5.1 Settings
Models CLIPScore $\uparrow$ DynamicDegree $\uparrow$ AestheticQuality $\uparrow$ FID $\downarrow$ FaceSim $\uparrow$
ID-Animator [8] 28.921 0.280 0.595 159.109 0.349
consisID [10] 30.610 0.871 0.594 200.400 0.414
EchoVideo 30.567 0.771 0.601 154.678 0.138
Table 1: Quantitative results. The best is in bold. Addressing the "copy-paste" issue that leads to reduced facial similarity.
Models CLIPScore $\uparrow$ DynamicDegree $\uparrow$ AestheticQuality $\uparrow$ FID $\downarrow$ FaceSim $\uparrow$
w/o IITF 31.418 0.722 0.579 254.469 0.025
w/o Two-Stage 29.480 0.606 0.587 155.557 0.135
EchoVideo 30.567 0.771 0.601 154.678 0.138
Table 2: Ablation study on IITF and training strategy. "w/o IITF" denotes the exclusion of IITF module. "w/o Two-Stage" indicates the omission of the second training stage.
Evaluation. To mitigate the influence of visual feature extractors on celebrity faces, we carefully selected 23 images of ordinary individuals from the FFHQ dataset [21] as test subjects, ensuring these individuals were not part of the training set. The selection aimed for balanced representation across gender, age, and ethnicity, while also considering diversity in facial poses, lighting conditions, and occlusions. We designed 23 prompts based on the methodology from VBench [51], corresponding to various scenarios, activities, and facial poses. The evaluation of the IPT2V model emphasized human-annotated results. We selected a baseline model and employed the Good Same Bad (GSB) metric for quantification. Evaluation metrics were categorized into two dimensions: video quality and identity preservation. The video quality dimension followed the established text-to-video evaluation framework, subdivided into sensory quality, adherence to instructions, physical simulation, and coverage quality. The identity preservation dimension involved annotators comparing which video maintained closer identity consistency with the provided inference face image. Additionally, we utilized various tools, including VBench [51], to assess video generation quality and facial recognition embedding similarity for evaluating identity preservation effectiveness.
@@ 5.2 Results
Quantitative Results. We present the quantitative results of various methods across five dimensions in Table 1. The evaluation of generated videos is conducted using the dynamic degree and aesthetic quality metrics from VBench [51] to assess motion amplitude and visual appeal, respectively. Additionally, we employ the CLIPScore [52] to measure adherence to textual instructions. The Fréchet Inception Distance (FID) [53] is utilized to quantify the visual quality of the generated videos. Finally, we calculate the average cosine similarity of facial features between the reference image and the video frame image using FaceSim [50]. In terms of video quality, EchoVideo demonstrates a superior advantage; however, it falls short of ConsisID [10] in the dimensions of identity preservation and video dynamism. The remaining dimensions show comparable performance to other methods.
5.3 User Studies
We conducted a user study on the 529 videos to evaluate human preferences between our proposed EchoVideo, and ConsisID [10], with the latter serving as the baseline model. Each video pair was assessed by five annotators, and the outcome was determined through a voting mechanism. In non-tied cases, the majority decision prevailed. For tied votes, we used a scoring system: "good" scored 1, "same" scored 0, and "bad" scored -1. The cumulative score classified the result as "good" (> 0), "same" (= 0), or "worse" (< 0). The statistical analysis of the metrics, differentiated by gender, is presented in Table 3, where it can be observed that the proposed EchoVideo performs slightly worse compared to the ConsisID [10]. Nevertheless, it is evident that our model achieves state-of-the-art performance in this task.
@@ 5.4 Ablation Studies
All Gender Man Woman
Overall Quality -11.15% -9.42% -13.04%
Table 3: User study between EchoVideo and state-of-the-art method[10]. Each column displays the net good rate (good%-bad%).

### v2

@@ Abstract
Recent advancements in video generation have significantly impacted various downstream applications, particularly in identity-preserving video generation (IPT2V). However, existing methods struggle with "copy-paste" artifacts and low similarity issues, primarily due to their reliance on low-level facial image information. This dependence can result in rigid facial appearances and artifacts reflecting irrelevant details. To address these challenges, we propose EchoVideo, which employs two key strategies: (1) an Identity Image-Text Fusion Module (IITF) that integrates high-level semantic features from text, capturing clean facial identity representations while discarding occlusions, poses, and lighting variations to avoid the introduction of artifacts; (2) a two-stage training strategy, incorporating a stochastic method in the second phase to randomly utilize shallow facial information. The objective is to balance the enhancements in fidelity provided by shallow features while mitigating excessive reliance on them. This strategy encourages the model to utilize high-level features during training, ultimately fostering a more robust representation of facial identities. EchoVideo effectively preserves facial identities and maintains full-body integrity. Extensive experiments demonstrate that it achieves excellent results in generating high-quality, controllability and fidelity videos. The code and model are available at: https://github.com/bytedance/EchoVideo.
@@ 2.2 Identity-preserving Video Generation
Identity-preserving generation aims to retain distinct identity attributes in generated images or videos. Building on the remarkable success of diffusion models in text-to-image synthesis, numerous methods [42, 43, 44, 45, 46] have emerged that leverage these models for identity-preserving image generation. However, achieving identity preservation at the video level introduces additional complexity, as it requires maintaining identity consistency across frames while ensuring the naturalness of facial movements. Among the U-Net-based approaches, MagicMe [7] relies on fine-tuning with reference images, while ID-Animator [8] introduces a face adapter to encode facial features, allowing for tune-free operation. However, both methods have limitations regarding facial similarity and overall video quality. In contrast, several open-source DiT-based methods, such as ConsisID [10] and Magic Mirror [9], leverage the robust CogVideoX [37] model as a foundation, integrating facial information through cross-attention mechanisms to achieve identity preservation. FantasyID [47] incorporates a 3D facial geometry prior to ensure plausible facial structures during video synthesis. Despite generating high-quality videos and maintaining strong facial similarity, these methods suffer from significant “copy-paste” and “semantic conflicts” artifacts. To address these challenges, we propose EchoVideo, which incorporates a multimodal feature fusion module named IITF. This model effectively decouples facial features, enabling the generation of more stable and coherent portrait videos.
@@ 5.1 Settings
Model Identity Average $\uparrow$ Identity Variation $\downarrow$ Inception Distance $\downarrow$ Dynamic Degree $\uparrow$
ID-Animator [8] 0.349 0.032 159.11 0.280
ConsisID [10] 0.414 0.094 200.40 0.871
pika [12] 0.329 0.091 268.35 0.954
Ours 0.516 0.075 176.53 0.955
Table 1: Quantitative comparison with SOTA IPT2V methods. The best is in bold; the second best is underlined.
Evaluation. To mitigate the influence of visual feature extractors on celebrity faces, we carefully select 23 images of ordinary individuals from the FFHQ dataset [21] as test subjects, ensuring that these individuals are not part of the training set. The selection aims for balanced representation across gender, age, and ethnicity, while also considering diversity in facial poses, lighting conditions, and occlusions. We design 23 prompts based on the methodology from VBench [52], corresponding to various scenarios, activities, and facial poses. The evaluation of the IPT2V model emphasizes human-annotated results. We select a baseline model and employ the Good Same Bad (GSB) metric for quantification. Evaluation metrics are categorized into two dimensions: video quality and identity preservation. The video quality dimension follows the established text-to-video evaluation framework, subdivided into sensory quality, adherence to instructions, physical simulation, and coverage quality; The identity preservation dimension involves annotators comparing which video maintains closer identity consistency with the provided inference face image. Additionally, we utilize various tools, including VBench [52], to assess video generation quality and facial recognition embedding similarity for evaluating the effectiveness of identity preservation.
@@ 5.2 Results
Quantitative Results. We present the quantitative results of various methods across five dimensions in Table 1. The evaluation of generated videos is conducted using the dynamic degree and aesthetic quality metrics from VBench [52] to assess motion amplitude and visual appeal, respectively. Additionally, the Fréchet Inception Distance [53] is utilized to quantify the visual quality of the generated videos in the facial region. For identity preservation, we calculate the mean and standard deviation of the cosine similarity [54] between the facial features of the reference image and the video frame. In terms of facial similarity and video dynamics, EchoVideo demonstrates a superior advantage, while the performance in other dimensions is comparable to other methods.
@@ 5.3 Ablation Studies
Model Identity Average $\uparrow$ Identity Variation $\downarrow$ Inception Distance $\downarrow$ Dynamic Degree $\uparrow$
w/o IITF 0.025 0.017 268.09 0.898
w/o Two-Stage 0.495 0.073 183.46 0.951
w/o Mask Loss 0.386 0.077 180.80 0.941
EchoVideo 0.516 0.075 176.53 0.955
Table 2: Ablation studies. “w/o IITF” denotes the exclusion of IITF module. “w/o Two-Stage” indicates the omission of the second training stage. “w/o Mask Loss” refers to the absence of the box-based face mask loss.
@@
Mask Loss. We evaluated the impact of using face detection bounding boxes as mask regions for loss calculation. The quantitative results in Table 2 demonstrate that the incorporation of strong supervision for facial regions significantly enhances the similarity and stability of face generation.